### v1

# Phase engineering of MoS2 monolayers: A pathway to enhanced lithium-polysulfide battery performance

Abstract
[…]
I Introduction
The global transition to low-carbon energy systems demands high-performance energy storage technologies to support electric transportation, grid-scale systems, and portable electronics.Gielen et al. (2019); Mahlia et al. (2014); Kalair et al. (2021) Lithium–polysulfur (Li–S) batteries are among the most promising alternatives to conventional lithium-ion batteries (LIBs), offering a theoretical energy density of approximately 2600 Wh kg-1 and a specific capacity of 1675 mAh g-1, values two to three times higher than those of LIBs.Ji et al. (2009); Li et al. (2018a) Sulfur is also abundant, low-cost, and environmentally benign, making Li–S chemistry attractive for next-generation applications.Manthiram (2020); Su and Manthiram (2012a)
Li–S batteries operate through a multistep redox mechanism in which elemental sulfur (S8) is progressively reduced to form lithium polysulfides (Li2Sn), including long-chain species (Li2S6, Li2S8) and short-chain intermediates (Li2S2, Li2S4), before ultimately yielding solid Li2S during discharge.Su and Manthiram (2012b) This sequence governs critical aspects of cell performance, such as Coulombic efficiency, self-discharge, and cycle life. Lithium polysulfides (LimSn) thus play a central role in determining the electrochemical behavior of Li–S cells. Figure 1(top) shows the DFT-optimized geometries of key intermediates. Modeling this full polysulfide manifold is essential to reproduce the shuttle effect and to evaluate surface-engineering strategies that suppress it. Li et al. (2018b); Song et al. (2020a)
Despite their theoretical advantages, Li–S batteries face three persistent challenges. First, long-chain Li2Sn species are soluble in typical electrolytes and tend to diffuse toward the anode, causing self-discharge and capacity fading. Second, sulfur and its reduced forms exhibit extremely low electronic conductivity (e.g., $<10^{-30}$ S cm-1 for S8), which hinders charge-transfer kinetics. Third, the large volumetric expansion (up to 80%) during full lithiation compromises structural integrity and limits cyclability.Song et al. (2020b)
To address these limitations, two-dimensional (2D) materials have been proposed as functional hosts that can immobilize polysulfides, catalyze redox reactions, and buffer volume changes.Shao et al. (2019); Glavin et al. (2020); Khossossi et al. (2020); Fan and Huang (2022) Among these, molybdenum disulfide (MoS2) stands out for its rich polymorphism and tunable properties. This work focuses on the semiconducting 2H phase and the metallic 1T phase, which exhibit complementary functionalities as cathode materials.Voiry et al. (2015); Wang et al. (2019); Ni (2024) The 2H phase can suppress parasitic reactions due to its low conductivity, while the 1T phase promotes redox reactions through enhanced charge transfer and catalytic activity. Importantly, the 2H $\leftrightarrow$ 1T phase transition is reversible and can be induced via chemical doping, strain, or thermal treatment.Zheng et al. (2021); Lukowski et al. (2013) This tunability enables the design of hybrid architectures that combine the stability of semiconducting regions with the electrochemical reactivity of metallic ones, enhancing both rate performance and cycling stability.Li et al. (2017); Zhao et al. (2023)
In this work, we use density functional theory (DFT) to investigate the interaction between LimSn clusters and MoS2 monolayers in 2H and 1T phases. We compute adsorption energies, charge transfer, structural rearrangements, and delithiation pathways across the polysulfide series. By comparing how each polymorph stabilizes short- and long-chain intermediates, we uncover phase-engineering principles that enable effective polysulfide immobilization without compromising redox kinetics. These insights offer atomistic-level guidance for the design of high-performance Li–S battery cathodes based on tunable 2D materials.
Figure 1: Schematic representation of various lithium sulfide (LimSn) clusters and monolayer MoS2 structures. The top panel illustrates the geometries of the LimSn molecules studied, while the bottom panel depicts the 2H-MoS2 (left) and 1T-MoS2 (right) phases. For the MoS2 monolayers, both top (view along the c-axis) and side (view along the b-axis) perspectives are presented. The shaded regions in the bottom panel indicate the supercells under investigation. Lithium (Li) is shown in blue, molybdenum (Mo) in purple, and sulfur (S) in yellow.
II Computational Details
[…]
To calculate thermodynamic quantities, we complement periodic DFT with gas-phase quantum chemical calculations using the ORCA package (version 5.0.4).Neese (2012, 2022) The M06-L meta-GGA functionalZhao and Truhlar (2008) is used with the def2-TZVP basis set, along with the def2/J auxiliary basis for Coulomb fitting. All simulations employ a fine numerical integration grid (DEFGRID3) and tight SCF convergence criteria with an energy threshold of $1.0\times 10^{-8}$ a.u. Vibrational frequencies are computed within the harmonic approximation to obtain zero-point energies and thermal corrections to the enthalpy and Gibbs free energy. This molecular approach efficiently evaluates charge transfer, vibrational modes, and thermodynamic properties without the complexity of periodic boundary conditions.
III Results and Discussion
Herein presents a comprehensive first-principles analysis of lithium polysulfide interactions with MoS2 monolayers in both 2H and 1T phases. We begin by characterizing the adsorption geometries and energetics across the LimSn series, then by evaluating charge transfer, electronic structure modifications, and delithiation pathways. Finally, we incorporate vibrational and solvation effects in a gas-phase model to assess the thermodynamic viability of adsorption under realistic battery operating conditions.
III.1 Pristine MoS2
We consider two metastable polymorphs of molybdenum disulfide (MoS2): the semiconducting 2H phase with a trigonal prismatic structure (space group $P6_{3}/mmc$) and the metallic 1T phase with a distorted octahedral geometry (space group $C2/m$).Voiry et al. (2015); Krishnan et al. (2019) Figure 1(bottom) shows the atomic structures of both phases. We construct supercells containing 16 Mo and 32 S atoms for each case to enable direct comparison.
For 2H-MoS2, we use a $4\times 4\times 1$ supercell with lattice constants $a=b=12.61$ and a layer thickness of approximately 3.14 Å. The Mo-S and Mo-Mo bond lengths are 2.40 Å and 3.15 Å, respectively, in agreement with experimental and computational data.Fu et al. (2013) In contrast, the 1T phase is modeled using a $2\times 4\times 1$ supercell with $a=12.96$ and $b=12.59$, and a layer thickness of 3.47 Å. This distorted phase exhibits two distinct Mo-S bond lengths (2.41 Å and 4.50 Å) and three types of Mo-Mo separations (2.75 Å, 3.15 Å, and 3.76 Å), reflecting its reduced symmetry. Energetically, 2H-MoS2 is more stable, with the 1T phase lying approximately 0.19 eV per atom higher in energy. Nonetheless, experimental studies show that 1T-MoS2 can be stabilized via chemical doping, strain engineering, or electrochemical control,Lukowski et al. (2013); González et al. (2024), making it a viable candidate for battery applications where metallic character is advantageous.
Figure 2: LimSn adsorption properties on MoS2 monolayers. Panel (a) shows the adsorption energy ($E_{\text{ads}}$) for the 2H and 1T phases, panel (b) shows the charge transfer ($\Delta Q$) from the LimSn molecules to the MoS2 monolayer (positive values indicate charge transfer to MoS2), and panel (c) illustrates the vertical distance ($h$) between the molecules and the surface. Blue squares represent the 2H-MoS2 phase, and red circles represent the 1T-MoS2 phase.
III.2 Adsorption Properties of LimSn Molecules
We evaluate the adsorption of LimSn molecules on both 2H and 1T MoS2 monolayers using three key descriptors: adsorption energy ($E_{\mathrm{ads}}$), charge transfer ($\Delta Q$), and vertical distance ($h$) between the adsorbate and the surface, as shown in Figure 2. The adsorption energy is defined as:
$E_{\mathrm{ads}}=E_{\mathrm{full}}-E_{\mathrm{Li}_{m}\mathrm{S}_{n}}-E_{\mathrm{MoS_{2}}},$ (1)
where $E_{\mathrm{full}}$ is the total energy of the LimSn molecule adsorbed on MoS2, $E_{\mathrm{Li}_{m}\mathrm{S}_{n}}$ is the energy of the isolated molecule (computed in a $30\times 30\times 30$ cell), and $E_{\mathrm{MoS_{2}}}$ is the energy of pristine MoS2. A negative $E_{\mathrm{ads}}$ indicates an exothermic and thermodynamically favorable interaction.
Figure 2(a) shows that 1T-MoS2 exhibits stronger adsorption energies than 2H-MoS2, especially for smaller polysulfides such as Li2S and Li2S2. In the context of Li-S batteries, this suggests that 1T-MoS2 more effectively anchors polysulfides, suppressing the shuttle effect and improving cycling stability and energy efficiency.Manthiram (2020); Su and Manthiram (2012a)
We define the charge transfer $\Delta Q$ as:
$\Delta Q=Q_{\mathrm{full}}-Q_{\mathrm{MoS_{2}}},$ (2)
where $Q_{\mathrm{full}}$ is the total charge of the MoS2 monolayer with adsorbed LimSn, and $Q_{\mathrm{MoS_{2}}}$ is the charge of the pristine monolayer. A positive $\Delta Q$ indicates electron donation from the adsorbate to the surface. Figure 2(b) confirms this behavior across all configurations, with 1T-MoS2 exhibiting higher $\Delta Q$ values than 2H-MoS2, particularly for intermediate species such as Li2S2 and Li2S4. This enhanced charge transfer strengthens the binding interaction and further suppresses polysulfide diffusion.Manthiram (2020)
Figure 3: The density of states (DOS) of 2H-MoS2 with adsorbed LimSn molecules, compared with the pristine 2H-MoS2 DOS. Different panels correspond to the DOS for 2H-MoS2 with LimSn molecules: S8 in (a), Li2S in (b), Li2S2 in (c), Li2S4 in (d), Li2S6 in (e), and Li2S8 in (f), respectively. Blue lines represent the DOS of the 2H-MoS2 system with an adsorbed molecule, while the pristine DOS is shown as a shaded background for comparison.
Figure 4: The density of states (DOS) of 1T-MoS2 with adsorbed LimSn molecules, compared with the pristine 1T-MoS2 DOS (shaded region). The panel distribution and labeling are the same as in Fig. 3.
Figure 2(c) presents the vertical distance $h$ from the average topmost sulfur atoms of MoS2 to the nearest atom in the LimSn cluster. The 1T phase consistently yields shorter adsorption heights than the 2H phase, especially for Li2S and Li2S2, reinforcing its role as a more effective anchoring substrate.Su and Manthiram (2012a)
Overall, the 1T-MoS2 phase exhibits superior performance compared to 2H-MoS2 in terms of adsorption strength, charge donation, and interfacial proximity, indicating a greater ability to anchor polysulfides and facilitate redox reactions. This behavior is consistent with prior studies reporting that metallic TMDs generally interact more strongly with polysulfide species than their semiconducting counterparts.Hojaji et al. (2022); Liu et al. (2021) However, this enhanced interaction can also pose potential limitations: overly strong binding may hinder the desorption of reaction intermediates, particularly during the late stages of the redox cycle, and reduce reversibility under dynamic conditions.Huang et al. (2018); Hojaji et al. (2022) These findings suggest that while 1T domains offer significant advantages for polysulfide immobilization, careful control of their coverage and local environment is necessary to avoid kinetic bottlenecks associated with overbinding. Nevertheless, the weaker interaction of 2H-MoS2 may offer advantages by enabling controlled polysulfide mobility, potentially facilitating redox reversibility while limiting over-anchoring. This tradeoff highlights the value of hybrid systems that combine both phases to optimize sulfur cathode performance in Li-S batteries.
Our DFT results are in agreement with previous experimental reports. For instance, oxygen-doped MoS2 with partial 1T character achieves a discharge capacity of 1015 mAh g-1 with 78.5% retention over 300 cycles,Ren et al. (2024) consistent with enhanced charge transfer in metallic domains. Similarly, 1T-MoS2/MnO2 composites promote a transition from liquid-solid to solid-solid Li2S formation, enabling more uniform nucleation,Qing et al. (2024) while MCNT/MoS2/S hybrids maintain capacities near 460 mAh g-1 after 300 cycles,Guo et al. (2020) although the influence of conductive additives complicates direct interpretation.
III.3 Electronic Properties
Figures 3 and 4 present the density of states (DOS) for 2H- and 1T-MoS2 monolayers, respectively, following the adsorption of various LimSn species. For comparison, each panel also displays the DOS of the pristine MoS2 monolayer (shaded area), allowing for a direct assessment of the electronic modifications induced by polysulfide adsorption.
In 2H-MoS2, the adsorption of S8 and LimSn species introduces localized states near the valence band maximum (VBM), as seen in Figures 3(a)-(e). For example, Li2S4 and Li2S6 produce gap states just above the VBM, indicating mild hybridization and limited charge transfer.González et al. (2024) Notably, the band gap remains largely preserved after adsorption, reflecting the weak interaction and localized nature of these induced states. Similar behavior has been reported in other semiconducting TMDs, such as WS2 and MoSe2, where the intrinsic gap is retained despite adsorbate-induced features.Su and Manthiram (2012a) From a device standpoint, preserving the band gap can help limit parasitic conduction and reduce polysulfide shuttling.
By contrast, 1T-MoS2 exhibits pronounced changes in its DOS upon adsorption. Even for S8 [Figure 4(a)], significant hybridization occurs, with enhanced density of states appearing below the Fermi level (EF). This effect intensifies with lithium-containing species: Li2S4 and Li2S6 induce substantial DOS redistribution near EF [Figures 4(d) and (e)], consistent with the stronger charge transfer and binding energies reported in Section III.2. These changes reflect the metallic character of 1T-MoS2 and its ability to accept electrons from LimSn, facilitating improved electronic coupling and redox activity. These distinct responses arise from the inherent electronic differences between the two polymorphs. The semiconducting 2H-MoS2 shows minimal perturbation upon adsorption, with only weakly localized states near the band edges. In contrast, the metallic 1T phase undergoes substantial electronic restructuring, indicating stronger adsorbate-substrate interactions. These findings agree with previous studies reporting enhanced electron accommodation and catalytic activity in metallic TMDs.Manthiram (2020)
From a practical perspective, combining both polymorphs offers a promising strategy: 1T-MoS2 provides high conductivity and abundant active sites for redox reactions, while 2H-MoS2 may help stabilize the structure and suppress undesired electron leakage. This concept of phase complementarity is increasingly adopted in Li-S cathode design, leveraging the synergy between conductive and semiconducting domains to achieve both efficient charge transport and effective sulfur confinement.Lukowski et al. (2013)
Figure 5: NEB calculation for the Li2S4 delithiation process on 2H- and 1T-MoS2 surfaces. The external markers indicate the energies of metastable configurations along the reaction pathway. Blue lines represent the 2H phase, and red lines represent the 1T phase. Energies are given in electron volts (eV) relative to the reactants (energy of isolated molecule plus pristine surface).
Figure 6: Li2S4 delithiation process on the 1T-MoS2 surface. The adsorption energies ($E_{\text{ads}}$) for each step are shown below the structures. The color scheme follows Fig. 1.
III.4 Delithiation Pathway
Figure 5 displays the minimum energy paths for the delithiation of Li2S4 adsorbed on 2H- and 1T-MoS2, computed using the nudged elastic band (NEB) method. In both cases, the initial adsorption of Li2S4 is exothermic, with the 1T surface exhibiting stronger binding ($E_{\mathrm{ads}}=-1.96$ eV) than the 2H phase ($E_{\mathrm{ads}}=-1.06$ eV). This enhanced interaction reflects the higher affinity of metallic transition metal dichalcogenides (TMDs) for polar sulfur species, consistent with prior experimental and theoretical reports.Manthiram (2020) To gain mechanistic insight into the ion release process, we analyze the delithiation pathway in two successive steps: the removal of the first Li+ ion (transition state $\text{TS}_{a}$) and the subsequent detachment of the second Li+ ion (transition state $\text{TS}_{b}$). The corresponding energy barriers and intermediate stabilities reveal sharp contrasts in the kinetics and thermodynamics of the two MoS2 polymorphs.
First Delithiation Step ($\text{TS}_{a}$):
The initial delithiation step proceeds as $\text{Li}_{2}\text{S}_{4}\rightarrow\text{LiS}_{4}+\text{Li}^{+}.$ This reaction is kinetically hindered on the semiconducting 2H surface, with an activation energy of 1.54 eV and the LiS4 intermediate positioned 1.40 eV above the initial state. Such a high barrier suggests limited ion mobility and sluggish redox kinetics under ambient conditions. In contrast, the same process on the 1T phase features a much lower barrier of 0.40 eV, and the intermediate lies only 0.20 eV above the starting configuration. These values indicate that Li+ detachment is significantly more accessible on 1T-MoS2, consistent with its superior electronic and ionic transport properties.Manthiram (2020)
Second Delithiation Step ($\text{TS}_{b}$):
The subsequent step releases the remaining lithium ion: $\text{LiS}_{4}\rightarrow\text{S}_{4}+2\text{Li}^{+}.$ Once again, the 1T surface enables a favorable pathway, with a barrier of 0.40 eV and a final S4 product that is thermodynamically stable ($E=-1.34$ eV relative to the initial Li2S4 state). In contrast, the same transformation on 2H-MoS2 is much less favorable, requiring an activation energy of 1.43 eV and producing an unstable S4 species 1.70 eV above the initial state. These results emphasize the dual kinetic and thermodynamic limitations of the 2H phase in enabling efficient polysulfide conversion.
Figure 6 complements the NEB results by depicting the fully relaxed atomic configurations and corresponding adsorption energies along the delithiation sequence on 1T-MoS2. Throughout the process-from Li2S4 to LiS4 to S4-the system maintains strong interfacial coupling, accompanied by a progressive increase in charge transfer ($\Delta Q$). Specifically, $\Delta Q$ increases from 0.25 e- in the initial state to 0.36 e- for LiS4, and further to 0.44 e- for the final S4 product. This monotonic growth in $\Delta Q$ reflects efficient charge accommodation by the metallic substrate, enabling stabilization of partially oxidized intermediates via delocalized electronic states. The smooth energy landscape and persistent binding interactions suggest a continuous and energetically favorable reaction path that supports rapid redox kinetics and effective sulfur utilization. These properties, combined with reduced reaction barriers, reinforce the role of 1T-MoS2 as a highly effective anchoring and catalytic platform for lithium polysulfide conversion in Li-S batteries.Manthiram (2020)
The energy landscape of the delithiation process differs markedly between the two MoS2 polymorphs. The 1T phase exhibits low and relatively uniform barriers between successive intermediates, indicative of a smooth reaction pathway with minimal structural rearrangement. In contrast, the 2H phase shows abrupt energy variations between states, indicating a slower and energetically demanding process.
Figure 6 illustrates the optimized structures and key electronic descriptors along the delithiation sequence on 1T-MoS2. The initial adsorption of Li2S4 is strongly exothermic ($E_{\mathrm{ads}}=-1.96$ eV), accompanied by a charge transfer of $\Delta Q=0.25\,e^{-}$. Upon removal of the first Li+ ion, the resulting LiS4 intermediate remains strongly bound ($E_{\mathrm{ads}}=-1.74$ eV), with an increased charge transfer of $\Delta Q=0.36\,e^{-}$. Following the second delithiation, the final S4 species also remains stable ($E_{\mathrm{ads}}=-1.34$ eV), while $\Delta Q$ rises to 0.44 e-. The progressive increase in $\Delta Q$ along the reaction coordinate reflects the ability of 1T-MoS2 to redistribute and accommodate charge efficiently, stabilizing partially oxidized intermediates. As reported in related metallic TMD systems,Lukowski et al. (2013) such charge delocalization correlates with enhanced redox kinetics and higher rate capability.
III.5 Comparison 1T-MoS2 Polymorph
[…]
Therefore, 1T-MoS2 offers a more balanced profile: it exhibits sufficient binding to suppress polysulfide migration while maintaining low kinetic barriers to sustain redox activity. This intermediate behavior avoids the desorption bottlenecks of overbinding observed in vacancy-rich 1T phases, suggesting that 1T may serve as a more versatile and tunable platform for Li–S battery cathodes.
III.6 Thermochemistry (Gas-phase approach)
To evaluate the thermodynamic viability of polysulfide adsorption, we compute Gibbs free energy differences, $\Delta G$, using a gas-phase cluster model. This approach enables the inclusion of enthalpic and entropic contributions without relying on full periodic vibrational calculations, which remain computationally demanding, particularly when combined with solvation models. This method enables efficient and internally consistent evaluation of vibrational properties and thermal corrections, yielding accurate free energy estimates at finite temperature while avoiding the complexity and potential artifacts of periodic phonon calculations under solvation. Moreover, cluster-based approaches have been shown to produce results in close agreement with periodic calculations for systems where adsorption effects are spatially localized, further validating their use in this context.
At the time of writing, solvation models are not natively implemented in VASP for vibrational analyses. While alternative schemes such as VASPsol are available, we adopt a quantum chemistry-based strategy successfully applied in prior studies.Reuter (2012); Kurfman et al. (2021); Scott et al. (2012) Specifically, we extract a representative molecular cluster from the 1T-MoS2 surface and place the LimSn species at its geometric center. The cluster is sufficiently large to suppress edge effects and retain the local structural and electronic environment of the periodic system.
The thermodynamic stability of adsorbed polysulfides depends not only on their binding strength at 0 K, but also on the temperature and the chemical environment-particularly the presence of solvent. These factors are critical for real operating conditions in Li-S batteries, where temperature variations and electrolyte composition influence desorption, shuttling, and redox reversibility. To capture these effects, we compute Gibbs free energy changes ($\Delta G$) for LimSn adsorption on 1T-MoS2 across a range of temperatures and under both gas-phase and solvated conditions. Within this framework, we compute the Gibbs free energy change for adsorption as:
$\Delta G=G_{\mathrm{full}}(P,T)-G_{\mathrm{Li}_{m}\mathrm{S}_{n}}(P,T)-G_{\mathrm{MoS_{2}}}(P,T),$ (3)
where each term corresponds to the free energy of the adsorbate-cluster complex, the isolated LimSn molecule, and the pristine MoS2 cluster, respectively.
We include solvation effects using the conductor-like polarizable continuum model (CPCM) to account for the electrolyte environment in operating Li–S batteries.Tomasi et al. (2005); Barone and Cossi (1998) As a representative medium, we adopt dimethyl carbonate (DMC), a widely used solvent in lithium-based electrolytes, characterized by a dielectric constant of $\varepsilon=3.12$ and a refractive index of $n_{D}=1.419$.Xu (2004) This implicit solvation model captures the dielectric screening provided by the electrolyte, stabilizing polar or charged species and modulating the entropic contribution to adsorption, thereby influencing the overall free energy landscape.
Temperature Without solvation With solvation
(K) $\Delta G$ (eV) $\Delta G$ (eV)
253.15 -0.461 -0.311
273.15 -0.419 -0.267
293.15 -0.377 -0.222
313.15 -0.335 -0.178
Table 1: Gibbs free energy change ($\Delta G$) for LimSn adsorption on 1T-MoS2 at different temperatures (1 atm), comparing gas-phase and solvated (CPCM-DMC) models.
Table 1 summarizes the computed Gibbs free energy changes ($\Delta G$) for LimSn adsorption on 1T-MoS2 across a range of temperatures, comparing gas-phase and solvated conditions. Adsorption remains thermodynamically favorable ($\Delta G<0$) in all cases; however, the presence of solvation consistently reduces the magnitude of $\Delta G$, a trend that becomes more pronounced at higher temperatures due to the increasing entropic cost of adsorption.
As shown in Table 1, implicit solvation leads to systematically less negative $\Delta G$ values, indicating that solvent screening weakens the adsorbate–surface interaction. This behavior aligns with prior reports showing that polar environments reduce electrostatic binding between charged species and surfaces.Andreussi et al. (2012); Tomasi et al. (2005) For instance, at 293 K, near typical battery operating temperatures, the free energy of adsorption decreases from $-0.377$ eV in the gas phase to $-0.222$ eV under solvation, representing a reduction of approximately 41%. This difference becomes even more significant at 313 K, where $\Delta G$ drops from $-0.335$ eV (gas) to $-0.178$ eV (solvated), reflecting the growing impact of entropic penalties at elevated temperatures.
From a design perspective, electrolyte formulation must balance two competing effects: interactions strong enough to immobilize polysulfides and suppress the shuttle effect yet sufficiently weak to avoid overbinding that could hinder redox kinetics. Although DMC solvation reduces adsorption strength, it may enhance molecular mobility and help prevent electrode passivation. Solvation also affects the temperature threshold at which adsorption becomes thermodynamically unfavorable. Our calculations place this crossover near 473 K in the gas phase and approximately 393 K under implicit solvation, showing the role of the solvent in tuning the thermal stability and operational regimes of adsorbed polysulfide species.Manthiram (2020); Su and Manthiram (2012a)
Figure 7: Infrared (IR) spectra of Li2S4, 1T-MoS2, and Li2S4 adsorbed on 1T-MoS2, calculated using ORCA-DFT. The upper panel corresponds to gas-phase calculations without solvation effects, while the lower panel includes solvation effects. The arrows indicate the main vibrational features induced by the Li2S4 molecule on the 1T-MoS2 surface.
III.7 Spectral Analysis:
Infrared (IR) spectroscopy provides a direct and experimentally accessible probe of molecular bonding environments and interfacial interactions. To complement the thermodynamic analysis, Figure 7 presents the calculated IR spectra of Li2S4 adsorbed on a 1T-MoS2 cluster, evaluated under both gas-phase and solvated conditions. These spectra, obtained via DFT calculations using ORCA, reveal substantial vibrational shifts upon adsorption, indicative of strong chemical coupling between the polysulfide and the metallic surface.
In the gas phase, Li2S4 exhibits pronounced absorption features between 100 and 500 cm-1, primarily attributed to Li–S stretching and bending modes. Compared to the isolated molecule, these modes experience moderate redshifts (10–20 cm-1), indicating bond softening and geometric rearrangement upon surface binding. This behavior agrees with previous findings that TMD substrates can modulate vibrational force constants and induce conformational distortions in adsorbed polysulfides.Su and Manthiram (2012a)
Further spectral modifications are observed when solvation is included via the CPCM model, particularly at lower wavenumbers. Both peak positions and intensities change, suggesting that the dielectric environment influences local electric fields and vibrational coupling. These effects are consistent with the thermodynamic trend of weakened adsorption in solution and indicate notable electronic and structural reorganization at the solid-liquid interface.Tomasi et al. (2005); Barone and Cossi (1998); Xu (2004)
Significantly, adsorption on 1T-MoS2 consistently induces downshifts in several Li–S modes, often associated with bond elongation or weakening due to interfacial charge transfer and substrate-induced strain.Zhao and Truhlar (2008) These vibrational signatures correlate with the increase in charge transfer ($\Delta Q$) and stronger adsorption energies ($E_{\mathrm{ads}}$), reinforcing the role of 1T-MoS2 as an electronically responsive and structurally adaptive host for polysulfide immobilization.
IV Final Remarks
This study employs first-principles calculations to evaluate the potential of two MoS2 polymorphs, 2H and 1T, as anchoring materials for lithium-sulfur (Li-S) battery cathodes. We assess their ability to suppress polysulfide shuttling and enhance redox kinetics through detailed analysis of their electronic structure, adsorption energetics, and thermochemical response. Our results demonstrate that the metallic 1T phase consistently outperforms the semiconducting 2H phase in adsorption strength, charge transfer, and electronic coupling. These properties enable stronger immobilization of lithium polysulfides and facilitate faster redox dynamics. Conversely, the 2H phase, despite its weaker interactions, may improve mechanical integrity and reduce parasitic conductivity, suggesting a complementary role in hybrid architectures. Thermodynamic calculations reveal that solvation substantially weakens adsorption, lowering the polysulfide desorption threshold by approximately 80 K. This trend underscores the importance of considering solvent effects when designing interfacial chemistries. Simulated infrared spectra further confirm strong chemical coupling between Li2S4 and the 1T surface, manifested as pronounced redshifts in vibrational modes, consistent with significant charge transfer and interfacial restructuring.
From a materials design perspective, our results suggest that optimizing the coverage of 1T domains can enhance polysulfide anchoring, but excessive metallic character may lead to overbinding and hinder desorptionHuang et al. (2018); Hojaji et al. (2022). Operational stability may be improved by maintaining temperatures below the calculated desorption crossover (393 K under solvation), while targeted phase or defect engineering, through doping, strain, or electrochemical control, can be employed to tune the 2H/1T ratio and activate interphase regions. Additionally, electrolyte formulation should reduce polysulfide solubility while preserving sufficient mobility to avoid passivation or irreversible binding.
This work highlights the promise of phase-engineered MoS2 as a modular platform for next-generation lithium-polysulfide cathodes. A synergistic approach integrating computational screening with targeted experimental efforts will be key to realizing durable, high-performance, and tunable energy storage systems.
Acknowledgments
JWG acknowledges financial support from ANID-FONDECY (Chile) grants N. 1220700 and 1221301. RAG acknowledges financial support from Basal Program for Centers of Excellence, Grant AFB220001 (CEDENNA)
Competing Interests
The Authors declare no Competing Financial or Non-Financial Interests.
Data Availability
The data that support the findings of this study are available from the corresponding author, upon reasonable request.